# arXiv:2603.00425v1[cs.LG] 28 Feb 2026

### 28 Feb 2026 v1

@@ F.3 Performance Standard Deviation
Llama-3.2-1B SFT 0.0 0.0 0.0 0.0
LoRA $8.0\times 10^{-1}$ $4.8$ $5.5\times 10^{-1}$ $7.4\times 10^{-1}$
@@
gemma-3-1b SFT 0.0 0.0 0.0 0.0
LoRA $6.1\times 10^{-1}$ $3.3$ $3.0\times 10^{-1}$ $9.3$
@@
Qwen 3 4B SFT 0.0 0.0 0.0 0.0
LoRA $3.3\times 10^{-1}$ $8.0\times 10^{-1}$ $7.2\times 10^{-1}$ $7.3\times 10^{-1}$
@@
Table 11 shows the standard deviation across the 5 runs of the numbers shown in Table 1. We report a single SFT run (deterministic), hence std dev is not applicable.
@@ F.4 Hardware Details
All models are trained using the AdamW optimizer Loshchilov2017DecoupledWD with a batch size of 8 (simulated via gradient accumulation where necessary for larger models). For SFT and LoRA, we employ a cosine learning rate scheduler with a warmup ratio of $0.1$ and $0.06$, respectively. In contrast, following the minimalist design of activation-based steering, our adapter and the ReFT baseline do not utilize weight decay, warmup, or a learning rate scheduler. BoolQ, Winogrande, and ListOps are trained for 1 epoch and GSM8K for 3 epochs.

### 06 Mar 2026 v2

@@ F.3 Performance Standard Deviation
Llama-3.2-1B LoRA $8.0\times 10^{-1}$ $4.8$ $5.5\times 10^{-1}$ $7.4\times 10^{-1}$
@@
gemma-3-1b LoRA $6.1\times 10^{-1}$ $3.3$ $3.0\times 10^{-1}$ $9.3$
@@
Qwen 3 4B LoRA $3.3\times 10^{-1}$ $8.0\times 10^{-1}$ $7.2\times 10^{-1}$ $7.3\times 10^{-1}$
@@
Table 12 shows the standard deviation across the 5 runs of the numbers shown in Table 1. We report a single SFT run, hence std dev is not applicable.
@@ F.4 Hardware Details
All models are trained using the AdamW optimizer Loshchilov2017DecoupledWD with a batch size of 8 (simulated via gradient accumulation where necessary for larger models). For SFT and LoRA, we employ a cosine learning rate scheduler with a warmup ratio of $0.1$ and $0.06$, respectively. In contrast, following the minimalist design of activation-based steering, our adapter and the ReFT baseline do not utilize weight decay, warmup, or a learning rate scheduler. BoolQ, Winogrande, ARC Challenge, AQuA and ListOps are trained for 1 epoch and GSM8K for 3 epochs.
Our RL experiments on section 6.2 uses 512 sequence length, 1025 completion length, batch size 6, gradient accumulation 4 steps and 6 rollout generations. We sweep 12 learning rates and average across 3 random seeds. We use DeepSeek’s default chat template for formatting.